6. A researcher studying a Parkinson-like disease in an isolated population discovered that several individuals who developed symptoms at an average age of 52.3 years have a D2S44 allele (R) that is common in this population but not found outside of the population. A study of members of the population and members of a similar population that does have the allele (+) was done and the numbers of individuals with Parkinson-like symptoms were recorded (Table).

Disease Normal 196 D2S44 allele + 307

a. Is there an association between the allele and the disease?

b. What evidence do you have to support your conclusion in part a?

c. Are the disease and the allele linked in this population?

d. A woman from this population has come to you for counseling. She has the allele but is too young (25) to show symptoms. What would you advise her about her future chances of developing the disease?
Answer #1

Answer:

A):

  • RRR resistance residual value is for M allele containing population is 145/(145+24)= 145/168=0.85 RRR Value.

B):

  • Yes marker and disease causing genes marker are linked in the population but the adult having these disease symptom attain after 59.5 years.

C):

  • If 25 year lady come for advice which has this gene of parkinsonism disease, then it is 100% sure that she will show symptom of parkinson disease after 58.5 years.
  • But we can suggest her for dopamine test at specific interval because this is the main enzyme which is found to be at low level during parkinson disease occur.
  • By doing test, she will know whether her body has parkinson symptoms or not. If symptoms appears then medicine should taken which will maintain dopamine level to prevent this disease.

                                              Please Rate My Answer.......Thank......u.....